# Probing Syntax in Large Language Models: Successes and Remaining Challenges

### v1

@@ Abstract
The syntactic structures of sentences can be readily read-out from the activations of large language models (LLMs). However, the “structural probes” that have been developed to reveal this phenomenon are typically evaluated on an indiscriminate set of sentences. Consequently, it remains unclear whether structural and/or statistical factors systematically affect these syntactic representations. To address this issue, we conduct an in-depth analysis of structural probes on three controlled benchmarks. Our results are three-fold. First, structural probes are biased by a superficial property: the closer two words are in a sentence, the more likely structural probes will consider them as syntactically linked. Second, structural probes are challenged by linguistic properties: they poorly represent deep syntactic structures, and get interfered by interacting nouns or ungrammatical verb forms. Third, structural probes do not appear to be affected by the predictability of individual words. Overall, this work sheds light on the current challenges faced by structural probes. Providing a benchmark made of controlled stimuli to better evaluate their performance.
@@ 1 Introduction
Understanding how word sequences are combined to form the meaning of a sentence is a central challenge to linguistics (Chomsky, 1957; 1995). Behavioral (Rayner, 1978; Frazier & Rayner, 1987; Gleitman, 1990) and neuroimaging research (Pallier et al., 2011; Dehaene & Cohen, 2011; Devauchelle et al., 2009; Caucheteux et al., 2021; Friederici, 2018; Santi & Grodzinsky, 2007) show that this phenomenon requires the brain to build latent structures that link the words of sentences. Critically, such “syntactic” structure are thought to be autonomous: the infamous sentence “colorless green ideas sleep furiously” shows that words can be syntactically linked even if each word is highly unpredictable, and thus lead to a meaningless sentence Croft (1995); Chomsky (1957). Yet, the neural and computational implementations of such autonomous syntactic structures remain largely unresolved (Grodzinsky & Friederici, 2006; Kaan & Swaab, 2002).
@@
Large Language Models (LLMs) now offer an unprecedented opportunity to revisit this historical question. Not only can these models generate coherent sentences, but their internal activations explicitly represent the syntactic structures long theorized by linguists (Belinkov et al., 2017; Tenney et al., 2019; Peters et al., 2018; Hale & Stanojević, 2024; Hewitt & Manning, 2019; Diego-Simon et al., 2024). Following earlier work on linear probing (Alain & Bengio, 2017; Conneau et al., 2018), the ‘Structural Probe’ (Hewitt & Manning, 2019) recently showed that LLMs spontaneously learn to build a subspace of activations, where the distance between two contextualized word embedding corresponds to the distance that separates two words in the syntactic tree. Diego-Simon et al. (2024) further showed that the angle between these two words represents the type of syntactic relations (e.g. ‘subject of’, ‘object of’).
@@
However, the performance of such Syntactic Probes has mostly been assessed with aggregated scores over large naturalistic corpora (Guarasci et al., 2021). Consequently, it is unclear whether the syntactic representations revealed by these LLM probes, identify autonomous representations of syntax (Croft, 1995; Chomsky, 1957), or whether they rather rely on other, heuristic or statistical factors.
@@
To address this question, we evaluate the performance of two state-of-the-art syntactic probes (Hewitt & Manning, 2019; Diego-Simon et al., 2024) under a variety of linguistic conditions. To this end, we leverage both naturalistic datasets and controlled stimuli designed to manipulate specific linguistic (linear distance, depth, number interference) and statistical features (word predictability). With these tools, we aim to (1) characterize the strengths and limitations of structural probes (2) compare them with human behavior, and ultimately (3) provide a benchmark to better evaluate structural probes.
@@ 2 Related work
Understanding syntax with linear structural probes has been an extensive area of research initiated with the seminal work of Hewitt & Manning (2019). These probes have been extended using spectral (Müller-Eberstein et al., 2022), polar (Diego-Simon et al., 2024), and orthogonal (Limisiewicz & Mareček, 2021) transformations providing insights about the underlying vectorial spaces hosting syntactic processing (Reif et al., 2019). This suite of probes assumes that linguistic structure can be linearly extracted from LLM representations. Such linearity of structure has recently been also been reported both throughout pre-training (Shai et al., 2024) and in-context learning (Park et al., 2025).
@@ 3.1 Data
The Structural (Hewitt & Manning, 2019) and Polar (Diego-Simon et al., 2024) Probes were trained on the Universal Dependencies English Web Treebank (UD-EWT) corpus  [^1] https://universaldependencies.org/treebanks/en_ewt/ (Silveira et al., 2014). UD-EWT contains 16,622 sentences. The corpus consists of 254,820 words and 16,622 sentences, extracted from online journals, blogs and informative websites. Every sentence in the corpus has been manually annotated following the Universal Dependencies (UD) formalism (Nivre et al., 2017).
@@ 3.2 Probes
$d_{T}(w_{i},w_{j})=\arg\min_{k}\left(\mathds{1}[\mathbf{A}^{k}_{i,j}>0]\right),$ (2)
@@ 3.3 Training
The Structural and Polar probes were trained on activations from layer 16 of Mistral-7B-v0.1 (Jiang et al., 2023), Llama-2-7b-hf (Touvron et al., 2023), and BERT-large (Devlin et al., 2019), as this layer has been shown to best encode dependency structures [^2] Mistral-7B-v0.1 and Llama-2-7b-hf are transformer-based LLMs with 32 layers and 7 billion parameters. BERT-large is a masked language model with 24 layers and 340 million parameters. (Diego-Simon et al., 2024).
@@ 4.1.1 Naturalistic data
We find that both linear distance and syntactic depth negatively impact probe accuracy, indicating that long-range and deeply-nested dependencies still challenge current probes. Remarkably, for very deep structure (depth=10), the probe’s accuracy is similar to what can be read out from the original space of activations of the LLM (fig. 2). By contrast, word surprisal shows little to no impact on probe accuracy. This results mean that the predictability of a word given its context does not affect the syntactic representation identified by the syntactic probes (fig. 2). Overall, these three results suggest that probe performance is more sensitive to structural properties of language than to statistical predictability.
@@
fig. 3 shows the signed feature importance for the different input variables. Overall, the results confirm that probe accuracy are most affected by the linear distance and the syntactic depth. By contrast, surprisals lead to weaker and less consistent impact fig. 2. These results reinforce our earlier findings, suggesting that probe performance is most impacted by both syntactic depth and linear distance.
@@ 4.1.2 Controlled sentences
Once again, our findings align with those from the UD-EWT dataset: linear distance and syntactic depth strongly impact probe accuracy. For linear distance, fig. 2 we find a linear distance of 3 words is enough to decrease probe accuracy by 0,4 points. A similar effect happens for syntactic depth fig. 2. Notably, for CE sentences, the ‘Activation Space’ baseline matches with the Structural and Polar probes. We argue that this is due to the subject and the verb being located at adjacent locations. Finally, for surprisal fig. 2, we observe opposing trends, as reflected in fig. 3, where accuracy slightly increases with head surprisal, whereas child surprisal exhibits the opposite effect. When comparing the Structural and Polar probes, we find that the Polar Probe exhibits a slower decay in accuracy with increasing linear distance fig. 2. We hypothesize that its more constrained objective, including additional syntactic information, fosters learning more robust syntactic representations, that rely less on surface heuristics.
@@
Figure 4: Probes share some similarities with human syntactic processing (Top left): Accuracy of the Polar Probe on the subject-verb dependency in grammatical PP sentences with varying nesting levels, comparing cases where the final attractor is congruent or incongruent with the subject. Congruent cases are harder for the probe. (Top right): Probability of different Polar Probe predictions in grammatical and ungrammatical PP sentences with one $P$. Predictions are grouped by which word—the attractor, the subject, both, or neither—binds to the verb. Because the attractor is linearly close to the verb, the probes are misled and produce mostly incorrect parses (red), outnumbering correct ones (green). In addition, ungrammatical cases increase the attraction effect. (Bottom left): Accuracy profile for Right Branching (RB) where the linear distance is constant across nestings, modifying syntactic depth. Probes show primacy and recency effects. (Bottom Right): Accuracy profile for Center Embedding (CE) sentences, inner dependencies, despite being deeper, are also closer together resulting in higher probe accuracies. Error bars represent the standard error in all plots.
@@ 4.2.1 Congruent nouns impact negatively on probe performance
More recent work (Lakretz et al., 2022; Lampinen, 2024; Marvin & Linzen, 2018) has shown that interfering nouns reinforce the confidence of LLMs when predicting verb forms during the next-token prediction task. (refer to Appendix A for more details regarding congruence and model prediction). We hypothesize that, in contrast, interfering nouns may cause confusion in the syntactic parsing task, which is not generative.
@@ 4.2.2 Ungrammatical verb forms impact negatively on probe performance
Analogous to the interfering effects observed in grammatical sentences, previous works (Marvin & Linzen, 2018; Hu et al., 2024; Ryu & Lewis, 2021) have shown that ungrammaticality also fosters attraction effects in LLMs when evaluated in a generative setting. Mirroring findings from psycholinguistic studies.
As illustrated in Figures 4 and B.3, we find that in PP sentences with varying levels of nesting, the Polar Probe exhibits a consistently higher error rate when encountering ungrammatical verb forms. However, we note that even in sentences with a single PP, the Polar Probe tends to bind the verb to both the subject and the attractor, resulting in a wrong parse.
@@
Remarkably, as shown in Section D.2 such failure is somewhat mitigated by the use of a masked language model, where attention heads operate bidirectionally.
@@ 4.2.3 Accuracy Profiles for Right Branching and Center Embedding
In contrast, accuracy in CE sentences improves with deeper nesting. This finding underscores the dominant influence of linear distance over syntactic depth as found in fig. 2.
5 Discussion and Limitations
@@
Notably, in relatively simple sentences with one Prepositional Phrase, the probe tends to bind both the subject and the attractor to the verb, resulting in a wrong parse. Such result puts forward a key challenge for linear probes, their sensitiveness to linear distance and syntactic depth makes them prompt to errors. To address these challenges, and better evaluate probes, we introduce a set of controlled sentences to serve as a benchmark. These carefully designed stimuli enable a linguistically motivated and systematic evaluation of probe performance.
@@
The failure cases may be attributed to the training distribution of structural probes, which is heavily skewed toward dependencies with a linear distance of 1. Alternatively, LLMs might encode syntactic trees on a non-linear manifold in hidden space, so linear probes would inevitably distort that structure. Lastly, the failures could be attributed to the underlying models, failing to faithfully encode syntactic structure.
As recently reported (Lampinen, 2024), prompting LLMs affects their downstream behaviour. In this work, we input the sentences one-by-one without any prior task context, an approach that may limit the quality of the LLMs’ internal representations. Modifying this setup to include contextualized prompts could yield linguistically richer latent representations and potentially enhance the performance of syntactic probes.
@@ 6 Acknowledgments
This project was provided with computer and storage resources by GENCI at IDRIS thanks to the grant 2023-AD011014766 on the supercomputer Jean Zay’s the V100 and A100 partition (PDS). This project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 945304 (PDS).

### v2

@@ Abstract
The syntactic structures of sentences can be readily read-out from the activations of large language models (LLMs). However, the “structural probes” that have been developed to reveal this phenomenon are typically evaluated on an indiscriminate set of sentences. Consequently, it remains unclear whether structural and/or statistical factors systematically affect these syntactic representations. To address this issue, we conduct an in-depth analysis of structural probes on three controlled benchmarks. Our results are fourfold. First, structural probes are biased by a superficial property: the closer two words are in a sentence, the more likely structural probes will consider them as syntactically linked. Second, structural probes are challenged by linguistic properties: they poorly represent deep syntactic structures, and get interfered by interacting nouns or ungrammatical verb forms. Third, structural probes do not appear to be affected by the LLMs’ predictability of individual words. Fourth, despite these challenges, structural probes still reveal syntactic links far more accurately than the linear baseline or the LLMs’ raw activation spaces. Taken together, this work sheds light on both the challenges and the successes of current structural probes and provides a benchmark made of controlled stimuli to better evaluate their performance.
@@ 1 Introduction
Understanding how word sequences are combined to form the meaning of a sentence is a central challenge to linguistics (Chomsky, 1957; 1995). Behavioral (Rayner, 1978; Frazier & Rayner, 1987; Gleitman, 1990) and neuroimaging research (Pallier et al., 2011; Dehaene & Cohen, 2011; Devauchelle et al., 2009; Caucheteux et al., 2021; Friederici, 2018; Santi & Grodzinsky, 2007) show that this phenomenon requires the brain to build latent structures that link the words of sentences. Critically, such “syntactic” system is thought to be autonomous: the infamous sentence “colorless green ideas sleep furiously” shows that words can be syntactically linked even if each word is highly unpredictable, and thus lead to a nonsensical sentence Croft (1995); Chomsky (1957). Yet, the neural and computational implementations of such autonomous syntactic structures remain largely unresolved (Grodzinsky & Friederici, 2006; Kaan & Swaab, 2002).
@@
Large Language Models (LLMs) now offer an unprecedented opportunity to revisit this historical question. Not only can these models generate coherent sentences, but their internal activations explicitly represent the syntactic structures long theorized by linguists (Belinkov et al., 2017; Tenney et al., 2019; Peters et al., 2018; Hale & Stanojević, 2024; Hewitt & Manning, 2019; Diego-Simon et al., 2024). Following earlier work on linear probing (Alain & Bengio, 2017; Conneau et al., 2018), the ‘Structural Probe’ (Hewitt & Manning, 2019) recently showed that LLMs spontaneously learn to build a subspace of activations, where the distance between two contextualized word embedding corresponds to the distance that separates two words in the syntactic tree. Diego-Simon et al. (2024) further showed that the angle between these two words can represent the type of syntactic relations (e.g. ‘subject of’, ‘object of’).
@@
However, the performance of such Syntactic Probes has mostly been assessed with aggregated scores over large naturalistic corpora (Guarasci et al., 2021). Consequently, it is unclear whether the syntactic representations revealed by these LLM probes identify abstract representations of syntax (Croft, 1995; Chomsky, 1957), or whether they rather rely on other, heuristic or statistical factors.
@@
To address this question, we evaluate the performance of two state-of-the-art syntactic probes (Hewitt & Manning, 2019; Diego-Simon et al., 2024) under a variety of linguistic conditions. To this end, we leverage both naturalistic datasets and controlled stimuli designed to manipulate specific linguistic (linear distance, depth, number interference) and statistical (word predictability) features. With these tests, we aim to (1) characterize the strengths and limitations of structural probes (2) compare them with human behavior, and ultimately (3) provide a benchmark to better evaluate structural probes.
@@ 2 Related work
Understanding syntax with linear structural probes has been an extensive area of research initiated with the seminal work of Hewitt & Manning (2019). These probes have been extended using spectral (Müller-Eberstein et al., 2022), polar (Diego-Simon et al., 2024), and orthogonal (Limisiewicz & Mareček, 2021) transformations providing insights about the underlying vectorial spaces hosting syntactic processing (Reif et al., 2019).
This suite of probes rests on the premise that linguistic structure is linearly readable from LLM representations. Recent work suggests that linear representations also encode belief states, both across pre-training (Shai et al., 2024) and during in-context learning (Park et al., 2025).
@@ 3.1 Data
The Structural (Hewitt & Manning, 2019) and Polar (Diego-Simon et al., 2024) Probes were trained on the Universal Dependencies English Web Treebank (UD-EWT) corpus ***https://universaldependencies.org/treebanks/en_ewt/ (Silveira et al., 2014). UD-EWT contains 16,622 sentences. The corpus consists of 254,820 words and 16,622 sentences, extracted from online journals, blogs and informative websites. Every sentence in the corpus has been manually annotated following the Universal Dependencies (UD) formalism (Nivre et al., 2017).
@@ 3.2 Probes
$d_{T}(w_{i},w_{j})=\arg\min_{k}\left(\mathds{1}[(\mathbf{A}_{i,j})^{k}>0]\right),$ (2)
@@ 3.3 Training
The Structural and Polar probes were trained on activations from layer 16 of Mistral-7B-v0.1 (Jiang et al., 2023), Llama-2-7b-hf (Touvron et al., 2023), and BERT-large (Devlin et al., 2019), as this layer has been shown to best encode dependency structuresMistral-7B-v0.1 and Llama-2-7b-hf are transformer-based LLMs with 32 layers and 7 billion parameters. BERT-large is a masked language model with 24 layers and 340 million parameters. (Diego-Simon et al., 2024).
@@ 4.1.1 Naturalistic data
We find that both linear distance and syntactic depth negatively impact probe accuracy, indicating that long-range and deeply-nested dependencies still challenge current probes. Remarkably, for very deep structure (depth=10), the probe’s accuracy is similar to what can be read out from the original space of activations of the LLM (fig. 2). By contrast, word surprisal shows little to no impact on probe accuracy. These results mean that the predictability of a word given its context does not affect the syntactic representation identified by the syntactic probes (fig. 2). Overall, these three results suggest that probe performance is more sensitive to structural properties of language than to statistical predictability. Suggesting a better performance generalization to nonsensical sentences rather than to sentences containing long-range or deep dependencies.
@@
fig. 3 shows the signed feature importance for the different input variables. Overall, the results confirm that probe accuracy are most affected by the linear distance and the syntactic depth. By contrast, surprisals lead to weaker and less consistent impact (fig. 2). These results reinforce our earlier findings, suggesting that probe performance is most impacted by both syntactic depth and linear distance.
@@ 4.1.2 Controlled sentences
Once again, our findings align with those from the UD-EWT dataset: linear distance and syntactic depth strongly impact probe accuracy. For linear distance, fig. 2 we find a linear distance of 3 words is enough to decrease probe accuracy by 0,4 points. A similar effect happens for syntactic depth fig. 2. Notably, for CE sentences, the ‘Activation Space’ baseline matches with the Structural and Polar probes. We suggest that this is due to the subject and the verb being located at adjacent locations. Finally, for surprisal (fig. 2), we observe opposing trends, as reflected in fig. 3, where accuracy slightly increases with head surprisal, whereas child surprisal exhibits the opposite effect. When comparing the Structural and Polar probes, we find that the Polar Probe exhibits a slower decay in accuracy with increasing linear distance fig. 2. We hypothesize that its more constrained objective, including additional syntactic information, fosters learning more robust syntactic representations, that rely less on surface heuristics.
@@
Figure 4: Probes share some similarities with human syntactic processing but fail on simple PP sentences. (Top left): Accuracy of the Polar Probe on the subject-verb dependency in grammatical PP sentences with varying nesting levels, comparing cases where the final attractor is congruent or incongruent with the subject. Congruent cases are harder for the probe. (Top right): Probability of different Polar Probe predictions in grammatical and ungrammatical with one $P$. Predictions are grouped by which word; the attractor, the subject, both, or neither—binds to the verb. Because the attractor adjacent to the verb, the probes are misled and produce mostly incorrect parses (red), outnumbering correct ones (green). Ungrammatical cases increase the attraction effect. (Bottom left): Accuracy profile for Right Branching (RB) where the linear distance is constant across nestings, modifying syntactic depth. Probes show primacy and recency effects. (Bottom Right): Accuracy profile for Center Embedding (CE) sentences, inner dependencies, despite being deeper, are also closer together resulting in higher probe accuracies. Error bars represent the standard error in all plots.
@@ 4.2.1 Congruent nouns impact negatively on probe performance
More recent work (Lakretz et al., 2022; Lampinen, 2024; Marvin & Linzen, 2018) has shown that congruent attractor nouns reinforce the confidence of LLMs when predicting verb forms during the next-token prediction task. (refer to Appendix A for more details regarding congruence and model prediction). We hypothesize that, in contrast, incongruent attractor nouns may cause confusion in the syntactic parsing task, which is not generative.
@@ 4.2.2 Ungrammatical verb forms impact negatively on probe performance
Analogous to the interfering effects observed in grammatical sentences, previous works (Marvin & Linzen, 2018; Hu et al., 2024; Ryu & Lewis, 2021) have shown that ungrammaticality also fosters attraction effects in LLMs when evaluated in a generative setting. Mirroring findings from psycholinguistic studies. As illustrated in Figures 4 and 13, we find that in PP sentences with varying levels of nesting, the Polar Probe exhibits a consistently higher error rate when encountering ungrammatical verb forms.
4.2.3 One PP suffices to elicit attraction effects
Remarkably, we note that even in simple sentences with a single PP, the Polar Probe tends to bind the verb to both the subject and the attractor, resulting in a wrong parse (Figures 4 and 13).
@@ 4.2.3 One PP suffices to elicit attraction effects
As shown in Figure 16 such failure is somewhat mitigated by the use of a masked language model, where attention heads operate bidirectionally.
@@ 4.2.4 Accuracy profiles for Right Branching and Center Embedding
In contrast, accuracy in CE sentences improves with deeper nesting. This finding underscores the dominant influence of linear distance over syntactic depth as shown in fig. 3.
5 Discussion, Limitations and Future work
@@
Notably, in relatively simple sentences with one Prepositional Phrase (eg: The keys to the cabinet are big.), the probe tends to bind both the subject and the attractor to the verb, yielding a wrong parse (Section 4.2.3).
Such result puts forward a key challenge for linear probes, their sensitiveness to linear distance and syntactic depth makes them prompt to errors. In the PP case, the probes’ sensitiveness to linear distance becomes especially apparent— the attractor being adjacent to the main verb is enough to confuse probes to yield a wrong syntactic parse.
To address these challenges, and better evaluate probes, we introduce a set of controlled sentences to serve as a benchmark. These carefully designed stimuli enable a linguistically motivated and systematic evaluation of probe performance.
@@
Several factors likely underlie the current challenges of structural probes. First, their training data is strongly skewed toward dependencies with a linear distance of one, introducing a bias into the learned linear probe during its training. Second, if LLMs encode syntax along a nonlinear manifold in hidden space, any strictly linear probe will inevitably distort that representation. Third, some errors may originate not from the probes themselves but from the LLMs, which may fail to encode syntactic structure with full fidelity. Lastly, prompting LLMs has been shown to affect their downstream behavior (Lampinen, 2024), contextualizing the sentences using a carefully designed prompt might result in linguistically richer representations, potentially enhancing the performance of syntactic probes.
We posit that exploring non-Euclidean probes that better align with the geometry of LLM representations is a promising direction for future work. Likewise, training syntactic probes on controlled stimuli to better capture long-range dependencies and deeper hierarchical structures remains an important area for continued research.
@@ 6 Acknowledgments
This project was provided with computer and storage resources by GENCI at IDRIS thanks to the grant 2023-AD011014766 on the supercomputer Jean Zay’s the V100 and A100 partition (PDS).
This project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 945304 (PDS).
This project received funding from PSL University under the grant agreement ANR-10-IDEX-0001-02 (EC).
This project received funding from the Département d’Études Cognitives (DEC) at ENS under the grant agreement FrontCog, ANR-17-EURE-0017 (EC).
This project received funding from the French National Research Agency (ANR) under the grant agreement ComCogMean, Projet-ANR-23-CE28-0016 (EC).